Joshua is a year older than his sister McKay. The product of their age is 33 more than 7 times Joshua’s age. How old are Joshua and McKay?
Answer by Edwin McCravy(20081)

McKay's age = M
Joshua's age = M + 1

       M(M + 1) = 7(M + 1) + 33
         M² + M = 7M + 7 + 33
         M² + M = 7M + 40
   M² - 6M - 40 = 0
(M - 10)(M + 4) = 0

M - 10 = 0;  M + 4 = 0
     M = 10;     M = -4  <-- ignore negative answer.

McKay is 10 and Joshua is 11.

Edwin